### v2

# These Are Not All the Features You Are Looking For: A Fundamental Bottleneck in Supervised Pretraining

Abstract
Transfer learning is a cornerstone of modern machine learning, promising a way to adapt models pretrained on a broad mix of data to new tasks with minimal new data. However, a significant challenge remains in ensuring that transferred features are sufficient to handle unseen datasets, amplified by the difficulty of quantifying whether two tasks are "related". To address these challenges, we evaluate model transfer from a pretraining mixture to each of its component tasks, assessing whether pretrained features can match the performance of task-specific direct training. We identify a fundamental limitation in deep learning models – an "information saturation bottleneck" – where networks fail to learn new features once they encode similar competing features during training. When restricted to learning only a subset of key features during pretraining, models will permanently lose critical features for transfer and perform inconsistently on data distributions, even components of the training mixture. Empirical evidence from published studies suggests that this phenomenon is pervasive in deep learning architectures – factors such as data distribution or ordering affect the features that current representation learning methods can learn over time. This study suggests that relying solely on large-scale networks may not be as effective as focusing on task-specific training, when available. We propose richer feature representations as a potential solution to better generalize across new datasets and, specifically, present existing methods alongside a novel approach, the initial steps towards addressing this challenge.
1 Introduction
Transfer learning has become a cornerstone of modern machine learning and artificial intelligence (Bottou, 2011; Collobert et al., 2011; Oquab et al., 2014; Bommasani et al., 2021). In its simplest form – supervised transfer learning, a relatively large deep learning network is pre-trained on a large and diverse dataset composed of cheap examples that are somehow related to the task of interest. In a second phase, the network is adapted using a much smaller dataset composed of examples that directly illustrate the task of interest. This second dataset is often too small to train a network of that size from scratch (Li et al., 2017). It is hoped that the representations constructed during the pretraining phase will be useful during the adaptation phase.
It remains an significant challenge to ensure that transferred features are sufficient to handle new, unseen datasets. A pretrained network that is missing crucial feature information may not perform on par with direct learning, that is, a network with all the features we could obtain by training directly on enough data to learn a target task.
Foundation model folklore suggests that, by pretraining these enormous models on "everything", they would learn "all the features (abilities) we could ever want." Should we invest into building ever larger all-purpose foundational models, or into collecting tasks-specific datasets for training smaller specialized models?
Positive findings can only validate transfer from a particular foundational model to a particular target task. We cannot say more because we do not know how to even state that the target task is related to the pre-trained model “abilities”.
Instead, we provide a negative result by exposing shortcomings that appear under very favorable conditions. We assume that the pre-training data is drawn from a mixture of distribution that includes the target task as a component, and we show that we cannot expect that an adapted model will always match a directly trained model.
If we cannot always match direct training when adapting a model pre-trained on a mix that includes the target task, we cannot expect it either when we know even less about the pre-training data.
We uncover an information saturation bottleneck that occurs when models cannot learn new features after they have encoded similar competing features. When networks are restricted to are restricted to learning only a subset of key features during pretraining, they lose crucial feature information that is essential to effectively adapt to new tasks, or even ones they have seen before. These lost features are nearly impossible to recover by simply fine-tuning on limited data.
Although this phenomenon was not previously recognized as such, we argue that it is pervasive in deep learning models. We analyze this phenomenon rigorously through a simple theoretical counterexample, and discuss several published papers whose empirical findings can be explained by a consistent loss of features that would be useful for transfer.
Additionally, factors such as the distribution imbalances or shuffling order of the training can determine which features a model can learn over time. Finally, we propose richer feature representations as a potential solution to improve generalization across new datasets and present existing methods alongside novel approaches which represent the beginning of addressing this challenge.
Section 2 presents a formal framework to assess effectiveness of transfer. Section 3 demonstrates through a simple example, that even in optimal conditions, a deep network might miss key features. Section 4 reveals a fundamental bottleneck of deep networks, supported by empirical evidence across existing studies. It also introduces "rich representations" as a potential solution. Section 5 presents a method of constructing rich representations that improves performance when transferred to unseen distributions, at no additional pretraining cost.
Why study Mixture Distributions?
This paper examines how real-world pretraining works by studying complicated datasets with many imbalanced subgroups. We use a mathematical model called mixture distribution to represent these complex datasets as a single cohesive structure made up of multiple probability distributions.
Mixture of Experts
This is distinct from the neural network architecture known as "mixture of experts" (MoEs), where specialized "expert" models learn different parts of a task. When training on real-world datasets with many components, each "expert" may focus on smaller number tasks, but still be bound by the same limitations.
2 Problem Statement
Consider a set of distributions $\{P^{[j]}(X,Y)\}_{j}$ representing data from different populations, and a mixture distribution $P^{[\mathrm{mix}]}(X,Y)$ formed by a weighted sum of these subdistributions. [^1] Note: This mixture is non-trivial, meaning that all component distributions are present.
$P^{[\mathrm{mix}]}(X,Y)=\sum_{j}\lambda_{j}\cdot P^{[j]}(X,Y)\quad\text{where}~{}\lambda_{j}>0,~{}\sum_{j}\lambda_{j}=1.$
Suppose that we pretrain a network on mixture $P^{[\mathrm{mix}]}$. Let $P^{[i]}$ denote a fixed arbitrary subdistribution. Assuming that we have enough data to directly train a network on $P^{[i]}$, we ask
Question: Do features pretrained using $P^{[\mathrm{mix}]}$ work nearly as well as features learned directly for target $P^{[i]}?$
2.1 Two Ways to Train
Consider a neural network $f$ of the following form
$f(X;\theta,\bm{\gamma})=\sum_{j}\,\underbrace{\varphi_{j}(X;\theta)}_{\text{feature extractor}}\cdot\underbrace{\gamma_{j}}_{\text{classifier}}$ (1)
This network has two main components.
The feature extraction layers ($\varphi_{j}(X;\theta)$) initially transform the raw input $X$ into real-valued features using parameters $\theta.$ Furthermore, these transformations may be non-linear, enabling the extraction of complex signals from the input data.
The linear classification layer ($\gamma_{j}$) combines linearly the extracted features to produce the final prediction. [^2] We explain in the following section why it is sufficient to study the linear case.
There are two ways to train such a model for distribution $P^{[i]}$ (Figure 1)
  1. 1.
    The model could be directly trained on the target distribution $P^{[i]}.$ In this case, the model learns both feature parameters $\theta$ and classifier weights $\bm{\gamma}$ optimized for the target distribution $P^{[i]}$. This yields the final trained model $f(\cdot\ ;\theta^{[i]},\bm{\gamma}^{[i]})$.
  2. 2.
    Alternatively, we could train a model through transfer learning (via linear probing).Footnote 2 The model is pretrained on a mixture distribution $P^{[\mathrm{mix}]}$. The feature parameters $\theta^{[\mathrm{mix}]}$ are frozen and only the linear classifier weights $\bm{\gamma}$ are fine-tuned on $P^{[i]}$. This is called linear probing.
The two approaches primarily differ in how they learn features. In direct training, a model is capable of capturing any representable feature from the target distribution $P^{[i]}$. However, in transfer learning, the model is restricted to selecting features based on a different pretraining task $P^{[i]}$, which includes $P^{[i]}$ as a component among many.
Under the constraints of transfer learning, it is unclear whether models pretrained on broad mixtures retain sufficient features to match the performance task-specific training.
To assess their effectiveness in preserving performance, we explore whether the direct training solutions $f(X;\theta^{[i]},\gamma^{[i]})$ of any component $P^{[i]}$ can be matched (or approximated) by linearly combining features $\varphi_{j}(X;\theta^{[\mathrm{mix}]})$ from pretraining on the mixture.
We assume access to unlimited training data, which is important for effective direct training. Theoretically, we ask whether the function space spanned by features $\varphi_{j}(X;\theta^{[\mathrm{mix}]})$ contains the direct training solutions $f(\dots,\theta^{[i]},\gamma^{[i]})$ trained on unlimited data.
Figure 1: Two ways to train a classifier for $P^{[i]}$: directly from $P^{[i]}$ or transferred from $P^{[\mathrm{mix}]}$.
2.1.1 Studying the Linear Case is Sufficient
Linear Probing:
We focus on the scenario of transfer learning via linear probing, where only the final classification layer is modified during fine-tuning. Thus, our results directly apply to linear probing.
General Fine-Tuning:
Our findings extend beyond linear probing to more common case where all network layers are adjusted. This can be written as a particular instance within our framework.
Consider the general case of fine-tuning a pre-trained neural network $f(X;\mathbf{w}^{[\mathrm{mix}]})$ that updates all of its weights for target $P^{[i]}$. Under the assumptions of fine-tuning, the final weights $\mathbf{w}^{[i]}$ remain close to the pre-trained weights $\mathbf{w}^{[\mathrm{mix}]}.$ Thus, $\mathbf{w}^{[i]}$ can be reasonably approximated as a first-order Taylor expansion around $\mathbf{w}^{[\mathrm{mix}]}$
[…]
Observe that $f(X;\mathbf{w}^{[i]})$ is a linear combination of ${\partial f}/{\partial w_{j}}(X;\mathbf{w}^{[\mathrm{mix}]})$, which are exactly the neural tangent kernel (NTK) features (Jacot et al., 2018). The fine-tuning operation effectively modifies only the linear coefficients $(w_{j}^{[i]}-w_{j}^{[\mathrm{mix}]})$. Thus, this is an instance of our abstract problem statement, in the space of these NTK features.
Non-Linear Classifier:
It is also common to fine-tune multiple non-linear classification layers of a network. The previous argument also applies when we modify some, not all, weights
2.1.2 Networks with Non-Zero Training Error
When training neural networks, it is common practice to optimize the model until the training error reaches zero. However, we cannot learn useful information from studying these.
Transfer has historically been used when there is insufficient data to directly learn a task. A network may achieve zero training error without generalizing perfectly to the entire distribution. When not all features have been learned, transfer to a subdistribution indicates how a network may generalize.
Our problem setup makes the assumption of unlimited training data. When a model has trained on everything, zero training error translates to zero test error. When a network already performs perfectly on all subdistributions, transfer to a component tells us nothing about how it may generalize.
2.1.3 Covariance
Proposition 2.1.
Suppose we find a feature $\varphi(X)$ correlated output $Y$ under the subdistribution $P^{[i]}$ Then, $\varphi(X)$ is correlated with $Y$ under any mixture $P^{[\mathrm{mix}]}=\sum_{j}\lambda_{j}P^{[j]}$, except maybe for a measure zero set of mixture coefficients $(\lambda_{j})$.
$\begin{split}\forall\neq 0\quad\implies\mathrm{Cor}_{P^{[\mathrm{mix}]};\lambda}(\varphi(X),Y)\neq 0\quad\text{for }\text{ a.e. }\mathbf{\lambda}\end{split}$
Proof.
Assume that this is not the case. Two variables are correlated if and only if they have nonzero covariance. Therefore, mixture coefficients $\lambda_{i}$ must satisfy the equation:
$0=\mathrm{Cov}^{[\mathrm{mix}]}(Y,\varphi(X))=\sum_{j}\lambda_{j}\mathrm{Cov}^{[j]}(Y,\varphi(X)).$
Under our assumption, at least one of the covariances $\text{Cov}^{[j]}$ is nonzero. Thus, this equation represents a measure zero subset of mixture coefficients. Outside of this subset, $\mathrm{Cov}^{[\mathrm{mix}]}(Y,\varphi(X))\neq 0$. ∎
Intuitively, the remaining components must exactly cancel out the positive covariance of $P^{[i]}$, which only occurs for precisely-balanced mixture coefficients.
This suggests that a feature interesting to a target $P^{[i]}$ (i.e. - correlated with Y) is also interesting to the pretraining distribution $P^{[\mathrm{mix}]}$. However, a nonzero covariance alone does not mean that this feature will be part of the solution $f(\theta^{[\mathrm{mix}]},\gamma^{[\mathrm{mix}]})$ of the pre-training problem (useful).
If the information $\varphi(X)$ provides about $Y$ has already been introduced by existing features in the solution, it is no longer useful. This is consequential when the training setup is biased towards learning sparse representations.
Sparsity in deep networks
Stochastic gradient learning algorithms in deep learning models often have an implicit sparsity bias (Gunasekar et al., 2017; Andriushchenko et al., 2023), a likely contributor to the effectiveness of deep learning. In an oversimplified account of the analysis of Andriushchenko et al., when using stochastic gradient descent with large step sizes, features are added and removed somewhat randomly, but those that contribute to reducing the training error tend to stick around. Depending on the random order in which features are discovered by the training algorithm, it may construct different sparse solutions.
Many practical regularization tricks such as early-stopping or weight decay have a similar effect. In particular, weight decay acts differently in the last layer of a deep network (spreading the weights over all the available features) than in the inner layers (quickly pruning features that do not rapidly help reducing the training error.)
3 Counterexample
This study explores whether the features learned from mixture distribution perform as well across component tasks $P^{[i]}$ as features learned specifically for each task. We introduce a simple counterexample demonstrating that this desired outcome does not hold.
Example Setup
Imagine a feature extractor which can learn only two feature extraction functions $\Phi(X)=(\varphi_{1}(X),\varphi_{2}(X))$. Depending on parameters $\theta,$ the space spanned by the selected features can have zero, one, or two dimensions.
Consider four individual distributions in this space, illustrated in Figure 3:
  • $P^{[1]}[\>\Phi(X){=}(+1,0),\>Y{=}+1\>]=\tfrac{1}{2}$ and $P^{[1]}[\>\Phi(X){=}(0,\pm 1),\>Y{=}-1\>]=\tfrac{1}{4}$.
  • $P^{[2]}[\>\Phi(X){=}(-1,0),\>Y{=}+1\>]=\tfrac{1}{2}$ and $P^{[2]}[\>\Phi(X){=}(0,\pm 1),\>Y{=}-1\>]=\tfrac{1}{4}$.
  • $P^{[3]}[\>\Phi(X){=}(0,+1),\>Y{=}-1\>]=\tfrac{1}{2}$ and $P^{[3]}[\>\Phi(X){=}(\pm 1,0),\>Y{=}+1\>]=\tfrac{1}{4}$.
  • $P^{[4]}[\>\Phi(X){=}(0,-1),\>Y{=}-1\>]=\tfrac{1}{2}$ and $P^{[4]}[\>\Phi(X){=}(\pm 1,0),\>Y{=}+1\>]=\tfrac{1}{4}$.
Each distribution is defined across three distinct points: two points with probability $1/4$, and the third point with probability $1/2$. Combining these forms a mixture $P^{[\mathrm{mix}]},$ supported over four points (Figure 3).
$P^{[\mathrm{mix}]}=\lambda_{1}\cdot P^{[1]}+\lambda_{2}\cdot P^{[2]}+\lambda_{3}\cdot P^{[3]}+\lambda_{4}\cdot P^{[4]}\quad\text{where}~{}\lambda_{i}>0,~{}\sum_{i=1}\lambda_{i}=1$
A distribution $P^{[\mathrm{mix}]}$ with non-zero mixture coefficients $\lambda_{i}$, therefore, offers positive examples at $\Phi(X)=(\pm 1,0)$ and negative examples at $\Phi(X)=(0,\pm 1)$.
The Optimal Classifier
Suppose that we optimize binary classifiers for each of the four distributions. The training error of each $P^{[i]}$ is minimized using one specific feature: $\varphi_{1}(X)$ for $P^{[1]}$ and $P^{[2]}$, $\varphi_{1}(X)$ for $P^{[3]}$ and $P^{[4]}$.
Now, consider a binary classifier optimized for $P^{[\mathrm{mix}]}.$ Since its four points cannot be linearly separated, the training error is minimized by misclassifying the least-weighted point and correctly classifying the other three. Depending on which point is ignored, the optimal classifier is one of the four solutions in Figure 3
Figure 2: Four subdistributions and a combined mixture distribution are represented by red points (labeled $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+1$) and blue points (labeled $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1$) Each subdistribution includes three points; two points contain an equal number of examples, while the third point contains twice as many. The size of each point reflects the number of examples it represents. The final mixture distribution is a weighted average of these four component distributions (not drawn to scale).
Figure 3: Four cases illustrate optimal classifiers for different mixture distributions. Each classifier uses a dotted line to separate red points (labeled $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+1$) from the blue points (labeled $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1$). Sincef points are not linearly separable, an optimal classifier ignores the least-weighted point (smallest) and classifies the remaining three points. Interestingly, this optimal classifier can be sparsely represented using just one feature, either $\varphi_{1}(X)$ or $\varphi_{2}(X),$ but cannot classify two of the four subdistributions.
Observe that the optimal classifier $P^{[\mathrm{mix}]}$ can be sparsely represented using just one of the two features, determined by the balance of the mixture. Hence, a deep neural network which exhibits sparsity bias will represent this optimal solution by learning just one feature. In other words, the network learns a feature space that does not contain the optimal solutions for two of its four subdistributions. Specifically, when it learns only $\varphi_{1}$, it cannot correctly classify $P^{[3]}$ and $P^{[4]}$; when it learns only $\varphi_{2}$, it cannot correctly classify $P^{[1]}$ and $P^{[2]}$.  [^3] Some mixtures have multiple clusters of equal minimal weight and, therefore, multiple optimal solutions. However, this is a set of negligible probability.
We find that training a deep network (with sparsity bias) on the mixture distribution yields a sparse representation which can classify only half of its subdistributions. Even in this advantageous setup, a model pretrained on a mixture may miss important features for its subtasks.
3.1 Assumptions
In this counterexample, we study the training error as an indicator of how well a model can adapt to its subdistributions. We make the following assumptions
Assumption 1: A Model Will Learn These Features(Papyan et al., 2020) finds that a neural network learns representations which "collapse" into the problem space features, $\varphi_{1}(X)$, $\varphi_{2}(X)$
Assumption 2: We Cannot Study Non-Linear Decision Boundaries A network with a non-linear decision boundary could learn both features and optimally classify all subdistributions. However, we are not interested in studying solutions with zero training error (see Section 2.1.2).
4 True in Practice?
Our simple counterexample illustrates how a model may struggle to learn crucial features during training, which creates biased performance on subtasks of the training distribution.
When an optimally-trained model can miss a crucial feature in such a simple setup, it suggests that deep networks may frequently miss important feature when trained on more complex datasets. In fact, the results of many already-published papers could be explained by this phenomenon. Instead of running redundant experiments, we describe this limitation and explore existing experiments below.
4.1 An Information Saturation Bottleneck
We identify a fundamental limitation of deep learning models that arises from their sparsity bias – an information saturation bottleneck. These networks struggle to learn new features when they have encoded similar competing features during training.
When networks are restricted to learning a subset of key features during pretraining, they permanently lose critical features needed for effective transfer, leading to inconsistent performance on data distributions–even components of the pretraining mixture.
The previous counterexample demonstrates that balance of a mixture may determine which features a deep network will learn over time. We provide empirical evidence for this and identify additional factors which influence the subset of features learned by a network.
4.2 Spurious Features
Table 1: Empirical Evidence (Left) Performance gap after recovering core features with transfer (Right) The simple supervised models trained by their supervised ’DASHA’ workflow outperform eight well Genomic Foundation Models (GFMs) on the diverse NT benchmark (Xu et al., 2025)
Trial Accuracy
Data Direct Transfer via LinProb Change
($99\%$ corr) $P^{[bal]}$ $P^{[\mathrm{mix}]}\to P^{[bal]}$
MNIST-Fashion $97\%$ $94\%$ ($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3\%$)
MNIST-CIFAR   $90\%$ $81\%$ ($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-9\%$)
Model Model Pretraining Average Average Mean Median
Size Base-Pairs Score $\uparrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
Enformer 252M 4B 0.569 11.86 27.73 27.91
NT-1000G (500M) 500M 20.5T 0.625 10.52 33.48 36.74
NT-1000G (2.5B) 2.5B 20.5T 0.656 7.0 36.58 40.86
NT-Multispecies (500M) 500M 174B 0.700 3.81 40.76 45.07
Foundation NT-Multispecies (2.5B) 2.5B 174B 0.697 4.08 40.51 45.52
Models DNABERT-2 117M 32.5B 0.680 6.88 38.65 43.59
HyenaDNA-1K 1.6M 3.2B 0.708 6.92 41.2 43.36
HyenaDNA-32K 1.6M 3.2B 0.630 10.22 33.96 36.93
Caduceus-PS 1.9M 35B 0.689 6.69 39.08 41.38
Caduceus-PH 1.9M 35B 0.725 4.69 42.63 45.01
Supervised Wide ResNet 2.0M 0 0.694 6.83 37.16 43.08
Models UNet 4.5M 0 0.68 7.78 38.67 42.69
DASHA (our workflow) 10.5M 0 0.761 3.69 46.33 49.08
Spurious features are superficial features which are correlated with but do not predict the label. For example, a bird classification model may incorrectly learn to use a water background to predict a waterbird and a land background to predict a land bird.
(Pezeshki et al., 2021) find that learning spurious features hinders the appearance of core features essential for accurate prediction. They find that spurious features to arise from bias towards certain classes in complicated mixtures of data, descriptive of real-world datasets. Core features cannot appear after the gradient is consumed by spurious features, which provide the same information but cannot be used for prediction. Class imbalances produce spurious features which "starve" the model of essential features.
(Kirichenko et al., 2023b) propose a promising mitigation to recover core features via transfer, by reweighting the remaining features without the spurious feature $\varphi_{\text{spu}}$. They consider that a model, pretrained on a biased mixture $P^{[\mathrm{mix}]}$, may still learn core features after learning easier spurious features. They remove $P\varphi_{\text{spu}}$ by transferring the network to a class-balanced subdistribution $P^{[bal]}$ (where $\mathrm{Cor}\big{(}\varphi_{\text{spu}}(X),Y\big{)}=0$). The remaining core features are optimized via linear probing.
Despite significant accuracy improvements, a performance gap $\{3\%,9\%\}$ remains between the adapted model and one trained directly on $P^{[bal]}$ (Table 1). Here, directly training on P[bal] produces a solution which includes all of the core features. In contrast, pre-training on $P^{[\mathrm{mix}]}$ gives a solution which does not. Linear probing on $P^{[bal]}$ cannot recover these features.
In this example of the bottleneck, pretraining on an imbalanced mixture produces misleading spurious features, which prevent the network from learning core features which are essential for prediction.
4.3 Genomic Foundation Models
(Xu et al., 2025) find that simple supervised CNNs can be easily trained to match the performance of transformer-based genomic foundation models (GFMs). Using their "DASHA" workflow, they train and evaluate supervised models on a diverse set of genomic tasks $P^{[\mathrm{adja}]}$ from the Nucleotide Transformer (NT) benchmark. They compare to the performance of well-known GFMs with up to 2.5B parameters, which are pretrained on a broad mixture $P^{[\mathrm{mix}]}$ of genomic tasks. They find the directly trained models to consistently outperform the GFMs across all aggregated metrics of $P^{[\mathrm{adja}]}$ benchmark (Table 1).
Certain GFMs such as Caduceus or NT-Multispecies(500M) perform almost as well as supervised networks on specific tasks of $P^{[\mathrm{adja}]}$, but miss features needed to perform well across all $P^{[\mathrm{adja}]}$. This suggests that large scale models learn some, but not all critical features for $P^{[\mathrm{adja}]}$ which appear in the extensive $P^{[\mathrm{mix}]}.$ While GFMs can excel in specific scenarios, their performance on new tasks is, ultimately, constrained by missing features that arise from information saturation.
If foundational models saturate regardless of scale, it may be more efficient invest resources into creating direct training data.
4.4 Rich Representations: The Beginnings of a Solution?
(Zhang and Bottou, 2023) propose learning richer representations to capture these missing features. They construct models by concatenating feature representations of individual models and demonstrate significant improvements in accuracy of transfer.
To construct rich representations, they concatenate four copies of ResNet50 trained separately on ImageNet1K, with different random seeds, into an ensemble model (ResNetCat4). They evaluate against a baseline model (ResNet50W2) of the same increased size, trained once on ImageNet1K.
Figure 4: Richer Representations via Concatenation An ensemble model (ResNet50Cat4) is created by concatenating four ResNet models trained separately on ImageNet using different random seed. It is evaluated against a baseline model (ResNet50W2), which is a single ResNet model of the same size, trained once on ImageNet.
Table 2: Supervised transfer learning from ImageNet to iNat18, Cifar100, and Cifar10 using linear probing. The cat4 ensemble concatenates representations of $4$ separately trained networks. The baseline uses a ResNet model of similar size is trained once. These models contain approximately four times the parameters of the original ResNet50.
Direct Transfer via LinProb
Method Arch Params ImageNet1k iNat18 Cifar100 Cifar10
original resnet50 23.5m 75.58 37.91 73.23 90.57
baseline resnet50w2 93.9m 77.58 37.34 72.65 90.86
ensemble cat4$\times$×resnet50 94m 78.15 46.55 78.19 93.09
They find the concatenated ensemble to perform comparably to the baseline on ImageNet1K, but far better when transferred to new datasets, ($+9\%$ on INat18) and ($+5.5\%$ on CIFAR100). Additional comparison to the original ResNet50 reveals that only increasing size of a model (such as the baseline) will worsen performance of transfer: (.$-.5\%$ on INat18), and (.$-.6\%$ on CIFAR100).
The performance gain of the concatenated ensemble can only be attributed to the features recovered by a richer, combined representation. This means that the four concatenated networks learn different subsets of features when trained on identical data shuffled differently.
This also explains why performance on ImageNet1K remains relatively unchanged. Observe that these subsets of features work similarly on ImageNet, that is, each set of features provides the same information to reduce training error. Combining these features marginally improves performance on ImageNet1K ($+2\%$). Hence, they are not learned during training.
Using a similar concatenation approach, (Zhang et al., 2022) demonstrate increased returns in accuracy for Self-Supervised Learning as scale increases from 100M to 1B parameters, as well as on Vision Transformers as scale increases from 100M to 400M parameters (Figure 4). In other words, these rich representations recover crucial features that otherwise would be lost and significantly increases performance accuracy without incurring additional training costs.
Figure 5: Increased Returns in Accuracy as Scale Increases (Left) SSL, 100M to 1B params: The concatenated methods (in red and purple) outperform the baseline (dotted blue curve) on both SWAV trained on unlabeled ImageNet1K (top), and SEER on INSTAGRAM1B (bottom). (Right) ViT, 100M to 400M params: Concatenated representations (purple) outperform the baseline (red) consistently during transfer in both original (top) and modified (bottom) vision transformers (ViT)
4.5 Connections to Multi-Calibration
Missing features fundamentally limits the ability of a model to generalize to new distributions outside the training distribution (OOD). (Wald et al., 2021) finds a link between OOD performance and multi-calibration, introduced by (Hebert-Johnson et al., 2018). Specifically, they argue that calibration leads to better OOD generalization, by creating invariant representations (Arjovsky et al., 2019). (Guy et al., 2025) successfully demonstrate this empirically for in-distribution generalization (IID). However, any remediation requires recovering these missing features.
5 Experiments
5.1 Rich Representations That Cost Nothing
We propose an analogous method to construct the rich representations, inspired by (Zhang et al., 2022), which concatenates networks pretrained for fewer steps. Unlike previous cat experiments that preserve model size (i.e. - space), our method preserves pretraining compute (time). Our findings reveal that the (time-cat) method improves generalization to new datasets while maintaining the cost of pretraining.
5.1.1 Method
Our approach concatenates feature representations from multiple ResNet50 pretrained on ImageNet1k for a shorter duration (Deng et al., 2009), (He et al., 2016).  [^4] Implementation can be found here: https://github.com/facebookresearch/richreps-timecat Each ResNet50 networks is trained identically except using different random seeds. We transfer these concatenated networks using linear probing on new datasets: iNaturalist18, CIFAR-100, CIFAR-10 (Van Horn et al., 2018), (Krizhevsky, 2009).
Our baseline uses a single ResNet50 model trained on ImageNet1k for 4̃50k steps, (reaching 0.65 accuracy on ImageNetv2). We compare how this baseline performs when transferred against several experiments which distribute total pretraining across a number of networks: two ResNet50 networks each trained for 200k steps, four ResNet50 networks each trained for 100k steps, and five ResNet50 networks each trained for 80k steps.
[…]
All ResNet50 models are pretrained using the MLCommons AlgoPerf training framework and the AdamWScheduleFree optimizer (Kasimbeg et al., 2025), (Defazio et al., 2024). The models are transferred using SGD.
5.1.2 Results
Our method yield considerable improvements in performance despite maintaining the same total pretraining computation. By combining more networks that have been pretrained for fewer steps, our experiments produce models which generalize significantly better to new datasets: ($+12.5\%$) on iNaturalist18 and ($+6\%$) on CIFAR-100.
As expected, adding more sets of features does not significantly improve performance ImageNet1K. Since ResNet50 models learn features that provide similar information to reduce training loss, combining them on the training data does not does not yield further improvements.
If concatenating more models trained on fewer steps can yield a significantly better performing model on new datasets, there may be considerable value left on the table.
Table 3: Transfer Performance of Concatenated Models, Each Trained For Fewer Iterations A baseline ResNet50 model (cat1) is trained for 450k steps on ImageNet1k. Each experimental method (catn) creates a models by concatenating multiple (n) ResNet50 networks, each trained independently for 400k/n steps. Overall computational effort of pretraining is maintained.
Pretraining Steps Transfer via LinProb
Method Per ResNet50 Total ImageNet1k Inat18 Cifar100 Cifar10
(BL) cat1 450k 450k 78.4% 33.7% 71.0% 90.4%
cat2 200k 400k 78.5% 40.4% 74.2% 91.2%
cat4 100k 400k 78.3% 42.9% $76.8\%$ 92.7%
cat5 80k 400k 77.8% 46.2% 77.1% 92.8%
6 Limitations
This paper identifies a behavior in deep neural networks, but does not claim to fully discover, solve, or explain this. First, this paper acknowledges prior work on related issues, such as spurious features from imbalanced datasets or poor generalization from foundation models. This paper builds on previous research by using their experimental results as a foundation for its own findings. While it presents rich representations as a potential solution, this paper also recognizes that this solution is incomplete. Instead, it serves as the initial steps towards understanding the problem. Finally, this paper focuses on specific toy examples and empirical evidence to provide insight into when and how this behavior might occur. These results are not comprehensive but provide an early phase of understanding, beginning with the classical case of fine-tuning.
7 Conclusion
This paper explores whether training models across a mixture of tasks provides the necessary features for each component. Our results indicate that supervised transfer methods, however useful, often permanently lose essential features without justification, potentially limiting their ability to generalize effectively. External factors like dataset imbalance or random seed limit what subpopulations a model can learn before training starts. Foundation models cannot alleviate this phenomenon with scale alone. We identify the beginnings of a solution to retrieve these lost features, revealing a fresh problem space with significant potential.
Appendix A Additional Empirical Evidence
This paper introduces a phenomenon called the information saturation bottleneck where models saturate information from encoded features and, consequently, miss key features. This can cause models to perform inconsistently across different subgroups of data.
This bottleneck appears not just in the cases discussed in this paper, but is widespread across the existing machine learning literature. Below are a few additional examples where this issue has been observed.
Biological Foundation Models
Beyond the work in this paper, a number of other works use supervised model to match the performance of larger-scale models.
  • Convolutional method can match the performance of Transformer-based foundation models [Yang et al., 2024]
  • Domain-specific generative model can outperform pretraining for single-cell biology tasks [Kedzierska et al., 2023]
Fairness
  • Fairness treatment of individuals within classification tasks [Dwork et al., 2012]
  • Identifying biased treatment (discrimination on sensitive attributes) in supervised learning [Hardt et al., 2016]
  • Incompatibility between fairness conditions and accuracy of risk scores [Kleinberg et al., 2016]
Robustness
  • Data augmentation for the ImageNet task [Balestriero et al., 2022], [Kirichenko et al., 2023a]
  • Group-DRO on water-bird, CelebA, MultiNLI tasks [Sagawa et al., 2020]